40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1344     Entry time: Mon Mar 2 03:57:44 2009
Author: Yoichi 
Type: Update 
Category: Locking 
Subject: Sunday night locking 
Tonight's locking started with a boot fest of the FE computers which were all red when I came in.
It also took me sometime to realize that C1:IOO-MC_F was returning always zero to tdsavg, causing the offloadMCF script to do nothing.
I fixed this by rebooting c1iovme and c1iool0.

Like Rob on the thursday night, I was only able to reach arm power around 10.
This time, I turned down the MC WFS gain to 0.02 (from 0.3).
I also checked gains of most of the loops (MICH, PRC, SRC, DARM, CARM-MCL, CARM-AO).
All the loops looked fine until the lock was lost suddenly. Also the spectrum of MC_F did not change as the arm power was ramped up.
Actually, I was able to reach arm power=10 only once because I spent a long time checking the loop gains and spectrum at fine steps of the arm power.
So it is quite possible that this loss of lock was just caused by a seismic kick.
ELOG V3.1.3-